OpenStudy (swissgirl):

Let m and n be integers. Then prove that m and n have different parity iff m^2-n^2 is odd.

OpenStudy (anonymous):

You can do this through contradiction, by first assuming m and n are both even, then assuming they're both odd, and showing neither can be true. Here's a hint: If m is odd, then m^2 is odd. If m is even, then m^2 is even.

OpenStudy (anonymous):

All squared even numbers are even and all squared odd numbers are odd. when you subtract an odd and a even the result is an odd number

OpenStudy (swissgirl):

noo i need to use the if and only if format

OpenStudy (anonymous):

You'll have to prove it backwards eventually, but I'd focus on proving it forwards first.

OpenStudy (swissgirl):

wait let me copy and paste the format

OpenStudy (anonymous):

\[ (2 p+1)^2-(2 q)^2=4 p^2+4 p-4 q^2+1 \]

OpenStudy (anonymous):

It is one line proof

OpenStudy (swissgirl):

wiat let me show u an example of the way it needs to be proved

OpenStudy (anonymous):

\[ (2 p)^2-(2 q+1)^2=4 p^2-4 q^2-4 q-1 \]

OpenStudy (anonymous):

\[ (2 p + 1)^2 - (2 q)^2 = 4 p^2 + 4 p - 4 q^2 + 1 = 2 (2 p^2 + 2 p - 2 q^2) + 1= 2 k +1 \] where \[ k=2 p^2 + 2 p - 2 q^2 \] So it is odd.

OpenStudy (anonymous):

@swissgirl did you get it?

OpenStudy (swissgirl):

yaaaa ok i was just looking at the format i gotta use

OpenStudy (swissgirl):

i gotta show that if p then q and if q then p

OpenStudy (swissgirl):

I thought i had to use a more complicating format

OpenStudy (swissgirl):

ok i gotcha then

OpenStudy (anonymous):

No, it is easy. Take one even 2p+1 and one even 2q and expand you get it in the from 2k +1 so it isodd.

OpenStudy (swissgirl):

that was great. thanks